Two cars A and B are moving along a straight road in the same direction with velocities of 25 km/h and 40 km/h, respectively. Find the velocity of car B relative to car A.


1
Expert's answer
2021-03-29T09:03:30-0400
"v_{B\ wrt\ A}=v_B-v_A=40\ \dfrac{km}{h}-25\ \dfrac{km}{h}=15\ \dfrac{km}{h}."